OpenStudy (anonymous):

write the slope-intercept form of the equation of the line through the given points. (3,4)(0,4)

OpenStudy (hoblos):

it is a horizontal line since both points have same ordinates so the equation is y=4

OpenStudy (anonymous):

thanks

Directrix (directrix):

(3,4) and (0,4) Use the point slope form of a line. slope = change in y over change in x = (4-4)/3-0) = 0 Pick a point: say, (0,4) y - 4 = 0 (x - 0) y - 4 = 0 y = 4.